Write and solve an equation to find the angle, its complement, and its supplement in this problem: The supplement of angle is 36 degrees less than three times its complement.
Answer by richwmiller(17219)

a+s=180
s=180-a
a+c=90
c=90-a
s=-36+3c
substitute 180-a for s
substitute 90-a for c
180-a=36+3*(90-a)
I leave the solving to you